# Fast Preemption: Forward-Backward Cascade Learning for Efficient and Transferable Proactive Adversarial Defense

### v5

@@ 5.3 Evaluation metrics
We used clean accuracy and robust accuracy for performance evaluation, where higher accuracy is better from the defense perspective. Here, accuracy refers to the ratio of predictions that match their ground-truth labels. Clean accuracy was computed for benign images (original and REs) while the robust accuracy was computed for AEs.
@@ 6 Performance
In this section, we present the end-to-end evaluation results and comparisons with benchmarks. The results were obtained in the settings listed in Tab. 3 of Appendix C. The time cost was recorded with an NVIDIA A100 40GB GPU, as described in Appendix E.
@@
Figure 4: Perturbations of adversarial attacks and proposed preemptive defenses. Observation of differences between original image and RE show that Fast Preemption perturbations are unnoticeable. Comparison of perturbations between AutoAttack and Fast Preemption show that preemptive defense is more likely to succeed when perturbations are similar.
@@ 6.1 Robustness against Adversarial Attacks
We evaluated transferable defense performance for the situation in which the defense has no knowledge of the AutoAttack [11] and victim models [6, 29]. Table 1 compares the performance of Fast Preemption with that of the benchmark defenses against AutoAttack [11]. Fast Preemption performance was better than that of the benchmarks in terms of clean and robust accuracies, whereas its time cost was competitive. For example, in comparison with the least time costly method WideResNet-94-16 [6] with the CIFAR-10 dataset, Fast Preemption had 1.6% and 11.9% better clean and robust accuracies (Standard $L_{\infty}$), respectively, with 0.04 more seconds per generation. Compared with the SOTA preemptive defense, Bi-Level [32], which had the second best accuracy, Fast Preemption had 2.5% and 5.1% better clean and robust accuracies (Standard $L_{\infty}$), respectively and was 700 times faster. Note that preemptive defenses can work together with training-time and test-time defenses. For example, when combined with a test-time defense, the performance of Fast Preemption was remarkably improved, as reflected in the last rows of CIFAR-10 and ImageNet in Tab. 1, respectively.
Figure 4 illustrates the results of qualitative analysis. A visual comparison of the original image and RE revealed that the perturbations generated by Fast Preemption are unnoticeable. Moreover, the “Success” and “Failure” cases indicate that preemptive defense is more likely to succeed when the perturbations are similar to the attack ones. Note that the perturbation illustration $\delta_{Gray}$ was drawn by first normalizing the true perturbations $\delta$ from $[-\epsilon,+\epsilon]$ to $[0,1]$ to produce $\delta_{RGB}$ and then transforming $\delta_{RGB}$ into grayscale using the National Television Standards Committee (NTSC) formula [30]:
@@ 6.2 Resistance against Reversion Attacks
We conducted reversion attacks to evaluate whether preemptive defenses are reversible. Apart from the proposed Preemptive Reversion attack, we used JPEG compression and adversarial purification because they require no knowledge of the defense. Preemptive Reversion was evaluated in both white- and black-box scenarios (see Sec. 4.2). The results are summarized in Tab. 2. The results shown in red approach the original accuracy, demonstrating that only our reversion attack was able to reverse the preemptive defense. The other reversion attacks failed because they further reduced clean accuracy (due to distortion). Nevertheless, Preemptive Reversion is only feasible in the white-box setting. Without full knowledge of the defense, Fast Preemption is irreversible. Figure 5 further supports our claim that the white-box reversion successfully erases most protective perturbations while the black-box reversion fails. However, the white-box reversion cannot fully remove the protective perturbations due to the randomization in Fast Preemption (i.e., randomly initialized FGSM).
@@
Moreover, the results in the fourth row of Tab. 2, which were obtained without using our evaluation protocol, demonstrate that our protocol for evaluating reversion attacks is more comprehensive. The results shown in the other rows demonstrate that our Preemptive Reversion in the white-box setting is the only feasible reversion attack as it is the only one for which clean accuracy was higher after reversion. However, the lower numbers in the fourth row indicate that the decrease in accuracy with Preemptive Reversion in the white-box setting was due to successful reversion, whereas the other decreases were due to distortion. Therefore, if only the results in the fourth row are considered, a decrease in accuracy does not necessarily indicate the effectiveness of a reversion attack.
@@ 7 Ablation Studies
We conducted ablation studies on the Fast Preemption framework and its learning algorithm using the CIFAR-10 dataset. We used AutoAttack (Standard) [11] against a black-box target model RaWideResNet-70-16 [35], employing $L_{\infty}$-norm perturbations with $\epsilon=8/255$. As demonstrated in Fig. 6, Fast Preemption significantly outperforms the SOTA Bi-Level method [32], increasing clean accuracy from 92% to 95.3%, improving robust accuracy from 80.1% to 85.6%, and reducing computation time from 28 seconds to just 0.04 seconds. These studies were conducted to identify the optimal configurations, with more extensive ablation results presented in Appendix F.
@@ 8 Conclusion
We introduce an efficient and transferable preemptive adversarial defense, termed Fast Preemption. This approach establishes a novel layer in the defense hierarchy, Proactive Defense (ours) → Training-Time Defense → Test-Time Defense, distinct from conventional test-time and training-time defenses. The proposed mechanism is plug-and-play, transferable, and requires no prior knowledge, empowering users to safeguard media autonomously without reliance on third-party interventions. Moreover, our method addresses the efficiency limitations of proactive defenses, achieving processing speeds 700 times faster than existing approaches, a critical improvement for real-world applications where substantial delays (e.g., 28 seconds per image [32]) are infeasible. Finally, our approach maintains or enhances clean accuracy while delivering state-of-the-art robust accuracy, overcoming the common trade-off observed in other defense methods that degrade clean accuracy.
Nevertheless, our defense shows limitations against the proposed reversion attack in the white-box setting. While we design the defense allowing user customizing the backbone model (serving as a user-side key) to counter reversion attacks, simplifying user requirements is a priority for future work. Further research should also explore the relationship between backbone models and transferability, as discussed in Sec. F.2.
@@
This supplementary material provides an in-depth discussion of individual related works (Appendix A), the algorithms underpinning the proposed adversarial defense and adaptive reversion attack (Appendix B), model architectures (Appendix D), and the experimental configurations for both attacks and defenses (Appendix C). Additionally, it details the implementation specifics (Appendix E) and presents comprehensive ablation study results Appendix F, covering aspects such as the classifier, backbone models, the Forward-Backward Cascade Learning algorithm, and the strategy of averaging multiple Fast Gradient Sign Method (FGSM) examples.
Appendix A Related Work
Appendix B Algorithms of Fast Preemption and
Preemptive Reversion
@@ Appendix B Algorithms of Fast Preemption and
Appendix C Experiment Settings
The full experimental settings used for evaluations are listed in Tab. 3.
Appendix D Model Architectures
We evaluated attacks and defenses using models with various architectures, as listed in Tab. 4.
@@ Appendix D Model Architectures
Appendix E Implementation
Our implementation partially leveraged the ART [33] and RobustBench [12] leaderboards by using PyTorch. The experimental machine was equipped with an NVIDIA A100 40GB GPU.
Appendix F Ablation Studies
Table 5: Performance with different classifiers.
       Ablation        Time Cost(s)        Transferable Defense
       Classifier        Accuracy(%)        Clean(%)        PGD(%)        AutoAttack(%)
       Ground Truths        100        28        97.2        86.9        83.1
       ResNeXt        95.9        28        95.4        85.7        82.7
       Bi-Level        89.1        29        91.9        83.9        81.0
F.1 Classifier
Table 6: Performance of different backbone models.
Backbone Robustness Clean (%) 89.1 93.9 88.1
PGD (%) 43.9 77.1 61.8
AutoAttack (%) 42.4 74.0 58.6
Time Cost (s) for Defense 28 69 26
White-Box Defense Clean (%) 95.5 95.8 93.8
PGD (%) 84.9 91.6 85.9
AutoAttack (%) 81.8 90.2 83.3
PGD (%) 85.7 89.7 87.6
AutoAttack (%) 82.7 88.0 85.9
Transferable Defense Victim: XCiT-L12 [13] Clean (%) 95.3 95.3 95.4
PGD (%) 72.7 78.6 79.6
AutoAttack (%) 70.8 75.7 77.3
F.2 Backbone Model
F.3 Forward-Backward Cascade Learning
Table 7: Performance for different $N$ and optimization algorithms.
       Ablation        Time Cost (s)        Transferable Defense
       Optimization        $N$ Samples        Clean (%)        PGD (%)        AutoAttack (%)
       PGD        1        0.7        95.3        87.5        85.8
       FGSM        1        0.04        95.4        86.8        85.4
       PGD        20        0.8        95.2        87.7        85.9
       FGSM        20        0.04        95.5        87.8        85.9
F.4 Optimization Algorithm
Appendix G Evaluation across More Models
Time Cost (s) N/A 28 0.35 0.04
RaWideResNet-70-16 [35] Clean 93.5 92.1 95.5 95.4
Robust 71.4 81.3 86.8 85.9
WideResNet-70-16 [50] Clean 93.3 91.5 95.5 95.3
Robust 71.2 80.9 86.4 86.0
WideResNet-34-10 [36] Clean 91.5 90.2 95.1 95.1
Robust 62.1 76.2 78.7 80.8
WideResNet-28-10 [52] Clean 93.4 91.5 95.4 95.6
Robust 64.8 77.7 81.3 83.2
PreActResNet-18 [51] Clean 84.2 87.3 90.1 91.6
Robust 42.4 54.7 55.4 61.2
ResNest152 [41] Clean 86.9 87.6 91.2 92.4
Robust 62.4 73.2 75.8 78.2
ResNet-50 [8] Clean 86.6 88.5 91.6 93.2
Robust 52.0 68.1 66.7 72.0
ResNet-18 [1] Clean 86.6 88.1 91.5 92.7
Robust 51.4 67.5 67.8 72.8
XCiT-L12 [13] Clean 92.8 91.2 95.1 95.4
Robust 57.9 71.3 74.6 76.0
Average Clean 89.9 89.8 93.4 94.1
Robust 59.5 72.3 74.8 77.3
  • Bold indicates the best performance.
Appendix H Defense against More Attacks
Our primary evaluation adheres to the RobustBench [12] protocol, employing the strongest attack, AutoAttack [11]. As detailed in Sec. 5.2, this includes tests with standard and randomized AutoAttack variants, encompassing white-box APGD [11], FAB [10], and black-box Square [4] attacks, alongside EOT settings [5].
In this section, we extend the evaluation to include three additional attacks: white-box Carlini-Wagner (CW) [7], black-box Deformation-Constrained Warping (DeCoW) [28], and black-box Meta Conditional Generator (MCG) [53], demonstrating the broad applicability and generalizability of our defense against diverse attack strategies.
As shown in Tab. 9, our defense achieves robust accuracy of 90.3%, 93.6%, and 87.3% against CW, DeCoW, and MCG attacks, respectively, outperforming SOTA adversarial training [6] and proactive defenses [32]. Notably, these accuracies exceed those obtained against AutoAttack, further highlighting the robustness and reliability of our primary evaluation.
      Threat Model       Attack       WideResNet-94-16 [6]       Bi-Level [32]       Fast Preemption (ours)
      White-Box       AutoAttack (Standard) [11]       74.0       82.1       86.3
      CW [7]       80.8       85.4       90.3
      Black-Box       DeCoW [28]       88.8       89.7       93.6
      MCG [53]       76.1       83.3       87.3
  • Bold indicates the best performance.
Appendix I Visual Quality
The noise budgets utilized (as detailed in Tab. 3) adhere to standard practices in adversarial machine learning, ensuring the preservation of visual quality and minimizing the likelihood of detection. As illustrated in Fig. 10, for 1,000 images from CIFAR-10 and ImageNet, Structural Similarity Index Measure (SSIM) [49] and Learned Perceptual Image Patch Similarity (LPIPS) [55] scores confirm that the visual quality of our robust examples (REs) is sufficiently high to serve as effective substitutes for the original images. Furthermore, our REs exhibit comparable (often superior) visual quality relative to adversarial examples (AEs) and outputs of existing defense methods.
I.1 Structural Similarity Index (SSIM)
where:
  • $\mu_{x}$ and $\mu_{y}$: Mean intensity values of $x$ and $y$, respectively.
  • $\sigma_{x}^{2}$ and $\sigma_{y}^{2}$: Variance of $x$ and $y$, respectively.
  • $\sigma_{xy}$: Covariance between $x$ and $y$.
    $C_{1}=(K_{1}L)^{2},\quad C_{2}=(K_{2}L)^{2},$ (16)
The SSIM value ranges from -1 to 1, where:
  • 1: Perfect similarity.
  • 0: No similarity.
  • Negative values indicate dissimilarity.
I.2 Learned Perceptual Image Patch Similarity (LPIPS)
$LPIPS(x,y)=$ (17)
where:
  • $l$: Layer index in the deep network.
  • $\phi_{l}(x)$: Feature map of $x$ at layer $l$ in the network.
  • $\hat{\phi}_{l}$: Normalized feature maps for each channel.
  • $H_{l}$, $W_{l}$: Height and width of the feature map at layer $l$.

### v6

@@ 5.1 Evaluation metrics
We evaluated performance using clean accuracy and robust accuracy, where higher values indicate better defense effectiveness. Clean accuracy was computed for benign images (original and REs), while robust accuracy was computed for AEs. Additionally, to assess image quality, we calculated the structural similarity index measure (SSIM) [49] and Learned Perceptual Image Patch Similarity (LPIPS) [55] relative to the original images. Higher SSIM and lower LPIPS values indicate subtler changes in the images.
@@
Threat Model Attack WideResNet-94-16 Bi-Level Ours
White-Box AutoAttack 74.0 82.1 86.5
CW 80.8 85.4 90.3
Black-Box DeCoW 88.8 89.7 93.6
MCG 76.1 83.3 87.3
  • AutoAttack is the standard version. Bold indicates the best performance.
@@ 6 Performance
The experiments were conducted using the settings outlined in Tab. 8 of Appendix E. Time costs were recorded on an NVIDIA A100 40GB GPU.
@@
Figure 4: Perturbations of adversarial attacks and the proposed preemptive defenses. Visual comparison between the original image and the RE demonstrates that Fast Preemption perturbations are imperceptible. Additionally, comparison of perturbations between AutoAttack and Fast Preemption reveals that preemptive defense is more likely to succeed when perturbations are similar.
@@ 6.1 Robustness against Adversarial Attacks
We evaluated transferable defense performance under the condition where the defense lacks knowledge of both the attacks and victim models [6, 29]. Table 1 and Tab. 2 compares Fast Preemption with benchmark defenses against four attacks. Fast Preemption outperformed the benchmarks in both clean and robust accuracies while maintaining competitive time costs. For instance, compared to the least time-costly method, WideResNet-94-16 [6] on CIFAR-10 under Standard $L_{\infty}$ AutoAttack, Fast Preemption achieved 1.6% and 11.9% higher clean and robust accuracies, respectively, with only a marginal increase of 0.04 seconds per generation. Compared to the SOTA preemptive defense Bi-Level [32], which had the second-best accuracy, Fast Preemption improved clean and robust accuracies by 2.5% and 5.1%, respectively, while being 700 times faster. Notably, preemptive defenses can be combined with training- and test-time defenses. For example, when paired with a test-time defense, Fast Preemption’s performance was significantly enhanced, as shown in the final rows of CIFAR-10 and ImageNet in Tab. 1.
Figure 4 provides a qualitative analysis, comparing the original images with their corresponding REs. The perturbations introduced by Fast Preemption are barely perceptible. The average SSIM and LPIPS scores for 2,000 examples from CIFAR-10 and ImageNet are 0.97 and 0.01, respectively, both surpassing those of standard AutoAttack and the SOTA preemptive defense Bi-Level [32] (more details in Appendix A). Additionally, the “Success” and “Failure” cases indicate that preemptive defense is more likely to succeed when the perturbations closely resemble those of the attack. The perturbation illustration $\delta_{Gray}$ was generated by normalizing the true perturbations $\delta$ from $[-\epsilon,+\epsilon]$ to $[0,1]$, converting $\delta_{RGB}$ to grayscale using the NTSC formula [30].
@@ 6.2 Resistance against Reversion Attacks
We conducted reversion attacks to assess the reversibility of preemptive defenses. In addition to the proposed Preemptive Reversion attack, we employed JPEG compression and adversarial purification, which do not require knowledge of the defense. Preemptive Reversion was tested in both white-box and black-box scenarios (see Sec. 4.2). The results, summarized in Tab. 3, reveal that only Preemptive Reversion, indicated in red, succeeded in reversing the preemptive defense, bringing the accuracy closer to the original. Other reversion attacks failed, further degrading clean accuracy due to distortion. Notably, Preemptive Reversion is only effective in the white-box setting, as Fast Preemption remains irreversible without full knowledge of the defense.
Figure 5 corroborates our findings, showing that the white-box reversion successfully eliminates most protective perturbations, whereas the black-box reversion fails. However, due to the randomization inherent in Fast Preemption (i.e., randomly initialized FGSM), the white-box reversion cannot entirely remove all protective perturbations. The fourth row in Tab. 3, which was obtained without our evaluation protocol, further highlights the comprehensiveness of our evaluation method. The other results demonstrate that Preemptive Reversion in the white-box setting is the only reversion attack that improves clean accuracy after reversion. The lower numbers in the fourth row suggest that the accuracy decrease observed with Preemptive Reversion in the white-box setting resulted from successful reversion, while other accuracy declines were due to distortion. Consequently, a decrease in accuracy, when solely considered from the fourth row, does not necessarily indicate the success of a reversion attack.
@@ 7 Ablation Studies
We conducted an ablation study on the Fast Preemption framework and its learning algorithm using the CIFAR-10 dataset and AutoAttack (Standard) [11] ($L_{\infty}$-norm $\epsilon=8/255$) against a black-box target model, RaWideResNet-70-16 [35]. As shown in Fig. 6, Fast Preemption significantly outperforms the SOTA Bi-Level method [32], increasing clean accuracy from 92% to 95.3%, improving robust accuracy from 80.1% to 85.6%, and reducing computation time from 28 seconds to 0.04 seconds. These studies were conducted to identify optimal configurations, with further ablation results provided in Appendix B.
@@ 8 Conclusion
We propose Fast Preemption, an efficient and transferable preemptive adversarial defense that introduces a new layer in the defense hierarchy: Proactive Defense (ours) → Training-Time Defense → Test-Time Defense. Unlike conventional training- and test-time defenses, Fast Preemption is plug-and-play, transferable, and requires no prior knowledge, enabling users to safeguard media independently without relying on third-party interventions. Our approach also addresses the efficiency limitations of existing proactive defenses, achieving processing speeds 700 times faster than prior methods, an essential improvement for real-world applications, where significant delays (e.g., 28 seconds per image [32]) are impractical. Moreover, Fast Preemption maintains or improves clean accuracy while achieving SOTA robust accuracy, overcoming the common trade-off where defenses degrade clean accuracy.
Despite its advantages, our defense exhibits limitations against the proposed reversion attack in the white-box setting. While allowing users to customize the backbone model as a user-side key mitigates reversion attacks, further work is needed to simplify user requirements. Future research should also investigate the relationship between backbone models and transferability, as discussed in Sec. B.2.
@@
This supplementary material provides a comprehensive analysis of visual quality (Appendix A) and presents an extensive ablation study (Appendix B), investigating critical components such as the classifier, backbone models, the Forward-Backward Cascade Learning algorithm, and the strategy of averaging multiple Fast Gradient Sign Method (FGSM) examples. Additionally, it includes evaluation results against a broader set of models (Appendix C), elaborates on the algorithms underlying the proposed adversarial defense and adaptive reversion attack (Appendix D), and details the experimental configurations for both attack and defense scenarios (Appendix E).
Appendix A Visual Quality
The noise budgets utilized (as detailed in Tab. 8) adhere to standard practices in adversarial machine learning, ensuring the preservation of visual quality and minimizing the likelihood of detection. As illustrated in Fig. 7, for 2,000 images from CIFAR-10 and ImageNet, Structural Similarity Index Measure (SSIM) [49] and Learned Perceptual Image Patch Similarity (LPIPS) [55] scores confirm that the visual quality of our robust examples (REs) is sufficiently high to serve as effective substitutes for the original images. Furthermore, our REs exhibit comparable (often superior) visual quality relative to adversarial examples (AEs) and outputs of existing defense methods.
A.1 Structural Similarity Index (SSIM)
where:
  • $\mu_{x}$ and $\mu_{y}$: Mean intensity values of $x$ and $y$, respectively.
  • $\sigma_{x}^{2}$ and $\sigma_{y}^{2}$: Variance of $x$ and $y$, respectively.
  • $\sigma_{xy}$: Covariance between $x$ and $y$.
    $C_{1}=(K_{1}L)^{2},\quad C_{2}=(K_{2}L)^{2},$ (16)
The SSIM value ranges from -1 to 1, where:
  • 1: Perfect similarity.
  • 0: No similarity.
  • Negative values indicate dissimilarity.
A.2 Learned Perceptual Image Patch Similarity (LPIPS)
$LPIPS(x,y)=$ (17)
where:
  • $l$: Layer index in the deep network.
  • $\phi_{l}(x)$: Feature map of $x$ at layer $l$ in the network.
  • $\hat{\phi}_{l}$: Normalized feature maps for each channel.
  • $H_{l}$, $W_{l}$: Height and width of the feature map at layer $l$.
Appendix B Ablation Studies
Table 4: Performance with different classifiers.
       Ablation        Time Cost(s)        Transferable Defense
       Classifier        Accuracy(%)        Clean(%)        PGD(%)        AutoAttack(%)
       Ground Truths        100        28        97.2        86.9        83.1
       ResNeXt        95.9        28        95.4        85.7        82.7
       Bi-Level        89.1        29        91.9        83.9        81.0
B.1 Classifier
Table 5: Performance of different backbone models.
Backbone Robustness Clean (%) 89.1 93.9 88.1
PGD (%) 43.9 77.1 61.8
AutoAttack (%) 42.4 74.0 58.6
Time Cost (s) for Defense 28 69 26
White-Box Defense Clean (%) 95.5 95.8 93.8
PGD (%) 84.9 91.6 85.9
AutoAttack (%) 81.8 90.2 83.3
PGD (%) 85.7 89.7 87.6
AutoAttack (%) 82.7 88.0 85.9
Transferable Defense Victim: XCiT-L12 [13] Clean (%) 95.3 95.3 95.4
PGD (%) 72.7 78.6 79.6
AutoAttack (%) 70.8 75.7 77.3
B.2 Backbone Model
B.3 Forward-Backward Cascade Learning
Table 6: Performance for different $N$ and optimization algorithms.
       Ablation        Time Cost (s)        Transferable Defense
       Optimization        $N$ Samples        Clean (%)        PGD (%)        AutoAttack (%)
       PGD        1        0.7        95.3        87.5        85.8
       FGSM        1        0.04        95.4        86.8        85.4
       PGD        20        0.8        95.2        87.7        85.9
       FGSM        20        0.04        95.5        87.8        85.9
B.4 Optimization Algorithm
Time Cost (s) N/A 28 0.35 0.04
RaWideResNet-70-16 [35] Clean 93.5 92.1 95.5 95.4
Robust 71.4 81.3 86.8 85.9
WideResNet-70-16 [50] Clean 93.3 91.5 95.5 95.3
Robust 71.2 80.9 86.4 86.0
WideResNet-34-10 [36] Clean 91.5 90.2 95.1 95.1
Robust 62.1 76.2 78.7 80.8
WideResNet-28-10 [52] Clean 93.4 91.5 95.4 95.6
Robust 64.8 77.7 81.3 83.2
PreActResNet-18 [51] Clean 84.2 87.3 90.1 91.6
Robust 42.4 54.7 55.4 61.2
ResNest152 [41] Clean 86.9 87.6 91.2 92.4
Robust 62.4 73.2 75.8 78.2
ResNet-50 [8] Clean 86.6 88.5 91.6 93.2
Robust 52.0 68.1 66.7 72.0
ResNet-18 [1] Clean 86.6 88.1 91.5 92.7
Robust 51.4 67.5 67.8 72.8
XCiT-L12 [13] Clean 92.8 91.2 95.1 95.4
Robust 57.9 71.3 74.6 76.0
Average Clean 89.9 89.8 93.4 94.1
Robust 59.5 72.3 74.8 77.3
  • Bold indicates the best performance.
Appendix C Evaluation across More Models
Appendix D Detailed Algorithms
@@
Appendix E Experiment Settings
The full experimental settings used for evaluations are listed in Tab. 8. We evaluated attacks and defenses using models with various architectures, as listed in Tab. 9. Our implementation partially leveraged the ART [33] and RobustBench [12] leaderboards by using PyTorch. The experimental machine was equipped with an NVIDIA A100 40GB GPU.
@@ Appendix E Experiment Settings
Appendix F Related Works